OpenStudy (anonymous):

Giving MEDAL. Express 0.64 as a fraction in its lowest terms. Please show how you got the answer.

jimthompson5910 (jim_thompson5910):

what decimal place is the 4 in? The tenths place? hundredths? thousandths? or some other decimal place?

OpenStudy (anonymous):

hundreths place?

jimthompson5910 (jim_thompson5910):

correct, the 6 is in the tenths place

jimthompson5910 (jim_thompson5910):

so 0.64 means 64 hundredths

jimthompson5910 (jim_thompson5910):

telling us that 0.64 = 64/100 the last step is to reduce that fraction

OpenStudy (anonymous):

that's the part i'm having trouble with

OpenStudy (anonymous):

Well .64 written as a fraction is 64/100 since there are 2 decimal places it's over a 100. Then you can divide to find the simplest form. 64/100 = 32/50 = 16/25 I divided by 2 each time So answer is 16/25

jimthompson5910 (jim_thompson5910):

what factors do 64 and 100 have in common?

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

divide each by 2: 64/2 = 32 and 100/2 = 50 so 64/100 = 32/50

jimthompson5910 (jim_thompson5910):

can we reduce 32/50 further?

OpenStudy (anonymous):

i don't think so

jimthompson5910 (jim_thompson5910):

we can actually

OpenStudy (anonymous):

how?

jimthompson5910 (jim_thompson5910):

32 and 50 have a number in common (that's not 1) as a factor

OpenStudy (anonymous):

i don't know what that number is

jimthompson5910 (jim_thompson5910):

since we're ignoring 1, how about 2?

jimthompson5910 (jim_thompson5910):

32/2 = ?? 50/2 = ??

OpenStudy (anonymous):

32 divided by 2 is 16...

jimthompson5910 (jim_thompson5910):

yep, and 50/2 = ?

OpenStudy (anonymous):

25

jimthompson5910 (jim_thompson5910):

so 64/100 = 32/50 = 16/25

jimthompson5910 (jim_thompson5910):

16 and 25 do not have any common factors between them (other than 1, but we're ignoring this trivial factor), so 16/25 cannot be reduced further

OpenStudy (anonymous):

oh, ok. I get it now. Thanks so much!

jimthompson5910 (jim_thompson5910):

you're welcome